What is the number of neutrons in 5 g of D[2 is at the base of D]0 (D is H^2[1 is at the base of H])
1
Expert's answer
2020-11-14T13:51:47-0500

Deuterium has 1 neutron, and Oxygen has 8 neutrons (Assuming the molecule D2O is made up of Oxygen-16 only).

Therefore, D2O has 10 neutrons per molecule.

18g of D2O = Avogadro's Number (6.0221421 x 1023) of molecules

5g of D2O = ?

"=\dfrac{(5g) x (6.022x10^23)}{18g} =" 1.67278x1023 molecules


1 molecule = 10 neutrons

1.67278x1023 molecules = ?

"=\dfrac{1.67278x1023 molecules x 10 neutrons}{1molecule} = 1.67278x10^24neutrons"


= 1.67278x1024 neutrons